### v1

@@ Abstract
This paper introduces a novel approach that combines synthetic control with triple difference to address violations of the parallel trends assumption. While synthetic control has been widely applied to improve causal estimates in difference-in-differences (DID) frameworks, its use in triple-difference models has been underexplored. By transforming triple difference into a DID structure, this paper extends the applicability of synthetic control to a triple-difference framework, enabling more robust estimates when parallel trends are violated across multiple dimensions. The empirical example focuses on China’s “4+7 Cities” Centralized Drug Procurement pilot program. Based on the proposed procedure for synthetic triple difference, I find that the program can promote pharmaceutical innovation in terms of the number of patent applications even based on the recommended clustered standard error. This method contributes to improving causal inference in policy evaluations and offers a valuable tool for researchers dealing with heterogeneous treatment effects across subgroups.
@@ 1 Introduction
In this paper, I aim to initiate a conversation with the literature by proposing an approach that combines synthetic control with DDD. The synthetic control method, originally designed to improve the counterfactual for DID settings (Abadie et al., 2010), offers a powerful tool for constructing better control groups for DDD as well. However, its application to DDD settings, where treatment effects may differ across subgroups, has been underexplored. By extending the synthetic control approach to a transformed version of DDD, I show how it can improve the validity of estimates when parallel trends are violated across multiple dimensions, including treatment status, time, and subgroup characteristics.
The contribution of this paper is two-fold. First, I contribute to the discussion of the synthetic control methodology in DDD settings, showing how transforming DDD into DID can help facilitate parallel trends in more complex empirical applications. Second, I provide an empirical example in which China’s “4+7 Cities” Centralized Drug Procurement pilot program may promote the number of patent applications, to provide a practical guide for this approach in a real-world policy evaluation; moreover, in this empirical exercise, I compare the performance of this synthetic DDD method to standard DDD and DID estimators, illustrating the effectiveness and highlighting the potential benefits of this combined approach in improving the robustness of causal inference.
The remainder of this paper is structured as follows. In Section 2, I present the theoretical foundations of the transformed DDD model. Section 3 explains how synthetic control can be adapted for use in a DDD context based on an empirical example, including a brief explanation of the dataset and the policy under study; the results of the empirical analysis are discussed, and the performances of standard DDD, DID, and the proposed synthetic DDD estimators are compared. Finally, Section 4 concludes with a discussion of the broader implications for empirical research and policy evaluation.
@@ 2.1 A simple illustration
$Y_{it}=\alpha_{0}+\alpha_{1}(\text{Treat}_{i}\times\text{Post}_{t})+\mu_{i}+\lambda_{t}+\epsilon_{it}$ (1)
@@
$=$ $\mathbb{E}[Y_{0,it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{0,it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]$
@@
$\begin{split}\alpha_{1}=&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right)\end{split}$ (2)
@@
$Z_{it}=\beta_{0}+\beta_{1}\text{Post}_{t}+u_{it}$ (3)
@@
$\hat{Z}_{it}=\hat{\beta}_{0}+\hat{\beta}_{1}\text{Post}_{t}+\hat{u}_{it}$ (7)
@@ 2.2 Converting DDD to DID
Let us first consider a standard triple-difference model without covariates as follows:
$Y_{it}=\alpha+\delta D_{it}\times G_{i}+\gamma_{1}D_{it}+\gamma_{2}G_{i}\times Post_{t}+\mu_{i}+\lambda_{t}+\varepsilon_{it}$ (9)
@@
For individuals in the subgroup of interest ($G_{i}=1$), we can define the transformed outcome variable as
$W_{it}=Y_{it}-\mathbb{E}[Y_{it}\mid G_{i}=0,\text{Treat}_{i}=j,t=t]$ (10)
@@
Thus, the sample counterpart of $W_{it}$ is:
$\hat{W}_{it}=Y_{it}-\hat{\bar{Y}}_{\text{Non-target},j,t}.$ (12)
@@
$\hat{W}_{it}=\hat{\theta}_{0}+\hat{\theta}_{1}D_{it}+\hat{\mu}_{i}+\hat{\lambda}_{t}+\hat{e}_{it}$ (13)
@@
$-\left.(\mathbb{E}[Y_{0,it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{0,it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right]$
@@
$\widetilde{W}_{it}=Y_{it}-\mathbb{E}[Y_{it}\mid G_{i}=0,\text{Treat}_{i}=j,t=t,X_{it}=x]$ (14)
@@
$\hat{Y}_{it}=X^{\prime}_{it}\hat{\beta}_{\text{Non-target},j,t}\text{ for $\text{Treat}_{i}=j=0,1$ and $t$}$ (15)
Thus, the sample counterpart of the adjusted outcome variable for $G_{i}=1$ is:
$\hat{\widetilde{W}}_{it}=Y_{it}-\bar{X}^{\prime}_{j,t}\hat{\beta}_{\text{Non-target},j,t}$ (16)
where $\bar{X}^{\prime}_{j,t}=\frac{1}{\sum_{i}1\{G_{i}=1,\text{Treat}_{i}=j\}}\sum_{i\in\{i|G_{i}=1,\text{Treat}_{i}=j\}}X_{it}\text{ for $j=0,1$}$. We then estimate the following regression for individuals with $G_{i}=1$:
$\hat{\widetilde{W}}_{it}=\hat{\widetilde{\theta}}_{0}+\hat{\widetilde{\theta}}_{1}D_{it}+\hat{\widetilde{\mu}}_{i}+\hat{\widetilde{\lambda}}_{t}+\hat{\widetilde{e}}_{it}$ (17)
with $\hat{\widetilde{\theta}}_{1}$ being a consistent estimator for the DDD effect with covariates, under the assumption that the relationship between $X_{it}$ and $Y_{it}$ is correctly specified and stable across groups.
After transforming the DDD estimator to a DID estimator by demeaning, we are ready to discuss combining DDD with synthetic control in practice.
@@ 3 Opening an Avenue for Synthetic DDD
As suggested by Olden and Møen (2022), all procedures for DID can be applied to a transformed DDD. This opens an avenue for us to combine synthetic control with triple difference based on Arkhangelsky et al. (2021). When should we apply this? How exactly can we do so in empirical research and how does it perform? In the following subsections, I answer these questions based on an empirical example.
@@ 3.1 When?
$\text{NumPatents}_{it}=\alpha_{0}+\alpha_{1}(\text{Treat}_{i}\times\text{Post}_{t})+X^{\prime}_{it}\alpha_{2}+\mu_{i}+\lambda_{t}+\epsilon_{it}$ (18)
@@
However, as shown by Figure 1, neither A61 nor non-A61 tracks exhibit parallel trends in the outcome variable, and the parallel trends are likely violated not only along the treatment dimension but also along the dimension of patent track. While synthetic DID can help create a more accurate counterfactual when parallel trends are violated along the treatment dimension, it might not capture complexities across different subgroups in this situation. Thus, my derivations above open an avenue for synthetic DDD, which ensures that the counterfactuals for both the treatment and the control groups are robust and account for these multi-dimensional violations of parallel trends. There can also be complex interactions between treatment, time, and subgroups that cannot be fully captured by synthetic DID alone. Thus, situations like these motivate us to combine synthetic control with DDD.
@@ 3.2 How?
$\begin{split}\text{NumPatents}_{it}=&\gamma_{0}+\gamma_{1}(\text{Treat}_{i}\times\text{Post}_{t}\times\text{Med}_{i})+\gamma_{2}(\text{Treat}_{i}\times\text{Post}_{t})+\\ &\gamma_{3}(\text{Post}_{t}\times\text{Med}_{i})+X^{\prime}_{it}\gamma_{4}+\mu_{i}+\lambda_{t}+\epsilon_{it}\end{split}$ (19)
$\begin{split}\text{DiffNumPatents}_{it}=&\theta_{0}+\theta_{1}(\text{Treat}_{i}\times\text{Post}_{t})+\mu_{i}+\lambda_{t}+e_{it}\end{split}$ (20)
@@
As shown in the table, when we do not include any covariates, the transformed DDD estimation gives exactly the same point estimate of the treatment effect as the standard DDD does, although standard errors are different. Since DID tends to over-reject the null hypothesis by having a smaller standard error (Bertrand et al., 2004), we recommend addressing this issue by using clustered robust standard errors. In Stata, the regular DDD can be estimated using command “reghdfe”, while the transformed DDD can be estimated using command “xtreg” or “reghdfe”.
@@
Figure 2: When combining synthetic control with transformed DDD
@@
Then, we can follow Arkhangelsky et al. (2021) to combine the transformed DDD with synthetic control using the Stata command “sdid”. From Figure 2, we can tell that parallel trends are much more apparent than those in Figure 1 after the synthetic control weighting. The point etimate suggests a treatment effect of 33.6 more patent applications in A61 track relative to non-A61 track annually. The standard error based on Arkhangelsky et al. (2021)’s placebo procedure is only about 5.9, leading to a P-value smaller than 0.001. We may conclude that centralized drug procurement can promote innovation activities in terms of the number of patent applications.

### v2

@@ Abstract
This paper discusses a practical approach that combines synthetic control with triple difference to address violations of the parallel trends assumption. By transforming triple difference into a DID structure, we can apply synthetic control to a triple-difference framework, enabling more robust estimates when parallel trends are violated across multiple dimensions. The proposed procedure is applied to a real-world dataset to illustrate when and how we should apply this practice, while cautions are presented afterwards. This method contributes to improving causal inference in policy evaluations and offers a valuable tool for researchers dealing with heterogeneous treatment effects across subgroups.
@@ 1 Introduction
In this paper, I aim to initiate a conversation with the literature by proposing an approach that combines synthetic control (SC) with DDD. The SC method, originally designed to improve the counterfactual for DID settings (Abadie et al., 2010), offers a powerful tool for constructing better control groups for DDD as well. However, its application to DDD settings, where treatment effects may differ across subgroups, has been underexplored. By extending the SC approach to a transformed version of DDD, I show how it can improve the validity of estimates when parallel trends are violated across multiple dimensions (including treatment status, time, and subgroup characteristics) due to the exponential growth of the outcome value. [^1] In such situation, a synthetic DDD estimator may be needed when taking the logarithm is not a good idea, e.g., if the outcome takes the value of zero for some units. A synthetic DDD estimator opens an avenue for causal inference when synthetic DID still performs poorly.
The contribution of this paper is two-fold. First, I contribute to the discussion of the SC methodology in DDD settings, showing how transforming DDD into DID can help facilitate parallel trends in more complex empirical applications. Second, I provide an empirical example to provide a practical guide for this approach in a real-world policy evaluation; moreover, in this empirical exercise, I compare the performance of this synthetic DDD method to standard DDD and DID estimators, illustrating the effectiveness and highlighting the potential benefits of this combined approach in improving the robustness of causal inference.
The remainder of this paper is structured as follows. In Section 2, I present the theoretical foundations of the transformed DDD model. Section 3 explains how SC can be adapted for use in a DDD context based on an empirical example, including a brief explanation of the dataset and the policy under study; the results of the empirical analysis are discussed, and the performances of standard DDD, DID, and the proposed synthetic DDD estimators are compared. In Section 4, I lay out additional notes and cautions for readers who intend to apply this procedure. Finally, Section 5 concludes with a discussion of the broader implications for empirical research and policy evaluation.
@@ 2.1 A simple illustration
$Y_{it}=\alpha_{0}+\alpha_{1}(\text{Treat}_{i}\times\text{Post}_{t})+\mu_{i}+\lambda_{t}+\epsilon_{it},$ (1)
@@
$=$ $\mathbb{E}[Y_{0,it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{0,it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0].$
@@
$\begin{split}\alpha_{1}=&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right).\end{split}$ (2)
@@
$Z_{it}=\beta_{0}+\beta_{1}\text{Post}_{t}+u_{it}\text{, if }\text{Treat}_{i}=1,$ (3)
@@
$\hat{Z}_{it}=\hat{\beta}_{0}+\hat{\beta}_{1}\text{Post}_{t}+\hat{u}_{it}\text{, if }\text{Treat}_{i}=1,$ (7)
@@ 2.2 Converting DDD to DID
Let us first consider a standard triple-difference model with fixed effects but without covariates as follows: [^2] Note that, this is a fully saturated model as in Olden and Møen (2022), since $Treat_{i}\times G_{i}$, $Treat_{i}$, and $G_{i}$ are absorbed by $\mu_{i}$, while $Post_{t}$ is absorbed by $\lambda_{i}$. One may also consider heterogeneous treatment effect by replacing $\delta$ by $\delta_{i}$
$Y_{it}=\alpha+\delta D_{it}\times G_{i}+\gamma_{1}D_{it}+\gamma_{2}G_{i}\times Post_{t}+\mu_{i}+\lambda_{t}+\varepsilon_{it},$ (9)
@@
For individuals in the subgroup of interest ($G_{i}=1$), we can define the transformed outcome variable in $\text{Treat}_{i}=j$ ($j=0,1$) as
$W_{itj}=Y_{it}-\mathbb{E}[Y_{it}\mid G_{i}=0,\text{Treat}_{i}=j,t=t],$ (10)
@@
Thus, the sample counterpart of $W_{itj}$ is:
$\hat{W}_{itj}=Y_{it}-\hat{\bar{Y}}_{\text{Non-target},j,t}.$ (12)
@@
$\hat{W}_{itj}=\hat{\theta}_{0}+\hat{\theta}_{1}D_{it}+\hat{\mu}_{i}+\hat{\lambda}_{t}+\hat{e}_{it}\text{, if }G_{i}=1,$ (13)
@@
$-\left.(\mathbb{E}[Y_{0,it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{0,it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right].$
@@
$W_{itjx}=Y_{it}-\mathbb{E}[Y_{it}\mid G_{i}=0,\text{Treat}_{i}=j,t=t,X_{it}=x].$ (14)
@@
$\hat{Y}_{it}=X^{\prime}_{it}\hat{\beta}_{\text{Non-target},j,t}\text{ for $\text{Treat}_{i}=j=0,1$ and $t$}.$ (15)
Thus, the sample counterpart of the adjusted outcome variable is:
$\hat{W}_{itjx}=Y_{it}-x^{\prime}\hat{\beta}_{\text{Non-target},j,t},$ (16)
where $x$ is a specific value of $X_{it}$. We then estimate the following regression for individuals with $G_{i}=1$:
$\hat{W}_{itjx}=\hat{\theta}_{0}+\hat{\theta}_{1}D_{it}+\hat{\mu}_{i}+\hat{\lambda}_{t}+\hat{e}_{it}\text{, if }G_{i}=1,$ (17)
with $\hat{\theta}_{1}$ being a consistent estimator for the DDD effect with covariates, under the assumption that the relationship between $X_{it}$ and $Y_{it}$ is correctly specified and stable across groups.
After transforming the DDD estimator to a DID estimator by demeaning, we are ready to discuss combining DDD with SC in practice, when the extended parallel trends assumption with covariates is violated.
@@ 3 Opening an Avenue for Synthetic DDD
As suggested by Olden and Møen (2022), all procedures for DID can be applied to a transformed DDD. This opens an avenue for us to combine SC with triple difference based on Arkhangelsky et al. (2021). When should we apply this? How exactly can we do so in empirical research and how does it perform? In the following subsections, I answer these questions based on an empirical example.
@@ 3.1 When?
$\text{NumPatents}_{it}=\alpha_{0}+\alpha_{1}(\text{Treat}_{i}\times\text{Post}_{t})+X^{\prime}_{it}\alpha_{2}+\mu_{i}+\lambda_{t}+\epsilon_{it},$ (18)
@@
However, as shown by Figure 1, neither A61 nor non-A61 tracks exhibit parallel trends in the outcome variable due to its exponential growth in value in the treated “4+7” cities, and the parallel trends are likely violated not only along the treatment dimension but also along the dimension of patent track. While synthetic DID can help create a more accurate counterfactual when parallel trends are violated along the treatment dimension, it might not capture complexities across different subgroups in this situation. As shown in Figure A1, pre-treatment trends are not as parallel as we may want based on synthetic DID. Thus, my derivations above open an avenue for synthetic DDD, which ensures that the counterfactuals for both the treatment and the control groups are robust and account for these multi-dimensional violations of parallel trends. There can also be complex interactions between treatment, time, and subgroups that cannot be fully captured by synthetic DID alone. Thus, situations like these motivate us to combine SC with DDD. [^3] A caution about applying the synthetic DDD procedure to count data is given in Section 4.4.
@@ 3.2 How?
$\begin{split}\text{NumPatents}_{it}=&\gamma_{0}+\gamma_{1}(\text{Treat}_{i}\times\text{Post}_{t}\times\text{Med}_{i})+\gamma_{2}(\text{Treat}_{i}\times\text{Post}_{t})+\\ &\gamma_{3}(\text{Post}_{t}\times\text{Med}_{i})+X^{\prime}_{it}\gamma_{4}+\mu_{i}+\lambda_{t}+\epsilon_{it},\end{split}$ (19)
$\begin{split}\text{DiffNumPatents}_{it}=&\theta_{0}+\theta_{1}(\text{Treat}_{i}\times\text{Post}_{t})+\mu_{i}+\lambda_{t}+e_{it},\end{split}$ (20)
@@
As shown in the table, when we do not include any covariates, the transformed DDD estimation gives exactly the same point estimate of the treatment effect as the standard DDD does, although standard errors are different. They differ due to several factors related to how each model handles residual variance and degrees of freedom. For instance, regular DDD involves a larger number of interaction terms, such as $\text{Treat}_{i}\times\text{Post}_{t}\times G_{i}$, which increases the complexity of the residual variance. This complexity can lead to a different calculation of the error variance, particularly if there is variation in the subgroup interaction terms. The transformed DDD simplifies the model by transforming the outcome into first differences. This transformation can alter the distribution of the residuals, potentially leading to a more efficient estimator with lower standard errors. The simplification reduces the number of parameters to estimate, as fewer interaction terms are involved, which can affect how the standard errors are computed.
Since DID tends to over-reject the null hypothesis by having a smaller standard error (Bertrand et al., 2004), one may use clustered robust standard errors to address this issue. In Stata, the regular DDD can be estimated using command “reghdfe”, while the transformed DDD can be estimated using command “xtreg” or “reghdfe”.
@@
Figure 2: When combining SC with transformed DDD
@@
Then, we can follow Arkhangelsky et al. (2021) to combine the transformed DDD with SC using the Stata command “sdid”. From Figure 2, we can tell that parallel trends are much more apparent than those in Figure 1 after the SC weighting. The point etimate suggests a treatment effect of 33.6 more patent applications in A61 track relative to non-A61 track annually. The standard error based on Arkhangelsky et al. (2021)’s placebo procedure is only about 5.9, leading to a P-value smaller than 0.001. We may conclude that centralized drug procurement can promote innovation activities in terms of the number of patent applications.
4 Additional Notes
This section presents a more formal description of the synthetic DDD estimator. I describe the estimator, the underlying assumptions, and derive its large sample distribution, emphasizing the use of the transformed outcome variable $W_{itjx}$.
4.1 The synthetic DDD estimator
I extend the synthetic DID framework to accommodate the triple-difference setting, capturing heterogeneous treatment effects across subgroups. The estimator relies on constructing a counterfactual outcome for the treated units that uses a weighted combination of control units and pre-treatment periods. The transformation we apply to the outcome is described in Equations (14) and (16). Let $\omega_{i}$ denote the unit-specific weights for the control group and $\lambda_{t}$ represent the time-specific weights for pre-treatment periods. The synthetic DDD (SDDD) estimator is then defined as:
$(\hat{\delta}^{\text{SDDD}},\hat{\alpha},\hat{\mu},\hat{\theta})=\arg\min_{\delta,\alpha,\mu,\theta}\left\{\sum_{i\in\{i\mid G_{i}=1\},t}\left(W_{itjx}-\alpha-\mu_{i}-\theta_{t}-D_{it}\delta\right)^{2}\hat{\omega}_{i}^{sdid}\hat{\lambda}_{t}^{sdid}\right\},$ (21)
where $\hat{\omega}_{i}^{sdid}$ and $\hat{\lambda}_{t}^{sdid}$ are chosen based on Arkhangelsky et al. (2021)’s procedure.
4.2 Key assumptions
To ensure that the proposed estimator provides consistent and unbiased estimates of the treatment effect, two important assumptions should be noted:
  1. 1.
    Large Panel Assumption: The estimator assumes a large panel, where both the number of cross-sectional units $N$ and the number of time periods $T$ grow large. This is necessary for consistency and asymptotic normality.
  2. 2.
    Normality of Error Terms: The error terms $\varepsilon_{it}$ are assumed to be normally distributed. This is important for deriving the large sample distribution.
4.3 Large sample distribution
Following the methodology of Arkhangelsky et al. (2021), one can derive the large sample distribution of the synthetic DDD estimator. Under the assumption that both $N\to\infty$ and $T\to\infty$, the synthetic DDD estimator converges to a normal distribution. Specifically, we have:
[…]
where $\delta$ is the true treatment effect and $\sigma^{2}$ is the asymptotic variance of the estimator.
The variance $\sigma^{2}$ is determined by the residual variance and the structure of the weights. The SC approach minimizes the differences between treated and control units in the pre-treatment period using weights.
4.4 Applicability to the empirical setting
In the empirical application above, I have only six time periods, which challenges the large sample assumption. In practice, finite sample sizes can affect the precision of the estimates. To account for this, one may use a block bootstrap procedure to estimate the variance in cases where the number of time periods is limited.
Another challenge the empirical exercise faces is the assumption that the error terms are normally distributed. However, one may argue that even though the number of patent applications is count data, when aggregated over time or across regions, the Central Limit Theorem implies that the distribution of these aggregated counts may approximate normality, making the assumption of normally distributed error terms reasonable in large samples.
Nevertheless, the empirical example provided earlier may only serve as a working illustration of when and how to apply the synthetic DDD strategy. Researchers should always be cautious about the underlying assumptions for an appropriate use of SC in a DDD setting.
4.5 Why don’t we combine SC with DDD directly?
One might ask: why don’t we create a SDDD estimator by directly combining SC with triple difference in the following form: [^4] Here I do not include covariates since we can simply incorporate adjustment for them by applying SDDD to the residuals $Y_{it}^{Res}=Y_{it}-X_{it}^{\prime}\hat{\beta}$ of the regression of $Y_{it}$ on $X_{it}$.
$(\hat{\delta}^{\text{sddd}},\hat{\alpha},\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\mu},\hat{\theta})=\arg\min_{\delta,\alpha,\gamma_{1},\gamma_{2}\mu,\theta}\left\{\sum_{i,t}\left(\begin{aligned} &Y_{it}-\alpha\\ &-\delta(\text{Treat}_{i}\times\text{Post}_{t}\times G_{i})\\ &-\gamma_{1}(\text{Treat}_{i}\times\text{Post}_{t})\\ &-\gamma_{2}(\text{Post}_{t}\times G_{i})\\ &-\mu_{i}-\theta_{t}\end{aligned}\right)^{2}\hat{\omega}_{i}^{sddd}\hat{\lambda}_{t}^{sddd}\right\}$ (23)
where $\hat{\omega}_{i}^{sddd}$ and $\hat{\lambda}_{t}^{sddd}$ should be chosen based on a new procedure.
At first glance, this approach seems intuitive. The SC method could be used to construct weights $\hat{\omega}_{i}^{sddd}$ and $\hat{\lambda}_{t}^{sddd}$ for both the cross-sectional units and the time periods, helping to create a more accurate counterfactual for treated units. The interaction terms $(\text{Treat}_{i}\times\text{Post}_{t}\times G_{i})$, $(\text{Treat}_{i}\times\text{Post}_{t})$, and $(\text{Post}_{t}\times G_{i})$ could account for the heterogeneous treatment effects across subgroups in the DDD setup. However, there are several challenges and considerations that may prevent this approach from being ideal, or even necessary, in practice:
  1. 1.
    Computational Complexity: Equation (23) involves much more complex computations compared to Equation (21). Estimating the weights $\hat{\omega}_{i}^{sddd}$ for each cross-sectional unit, $\hat{\lambda}_{t}^{sddd}$ for each time period, and the parameters $\hat{\delta}^{\text{sddd}}$, $\hat{\alpha}$, $\hat{\gamma}_{1}$, $\hat{\gamma}_{2}$, $\hat{\mu}$, and $\hat{\theta}$ simultaneously would involve a much larger optimization problem. The increased dimensionality could lead to issues with overfitting, unstable weights, and increased sensitivity to the choice of covariates or model specification.
  2. 2.
    Interpretability: The interpretation of the weights becomes more complex when combining SC directly with DDD. In Equation (21), the weights represent how much each control unit or time period contributes to constructing the counterfactual for treated units in a DID setting. However, in Equation (23), treatment effects are heterogeneous across both subgroups and time periods. Constructing appropriate weights that reflect these multi-dimensional interactions becomes challenging. It may be difficult to determine whether the weights provide meaningful interpretations for policy analysis and whether they accurately capture the multi-dimensional heterogeneity that DDD models aim to address.
  3. 3.
    Potential Redundancy: The transformed DDD approach discussed earlier already simplifies the DDD estimator by reducing it to a DID structure. By transforming the outcome variable to account for differences between treated and control subgroups, we can apply the synthetic DID framework to the transformed outcome without explicitly constructing a new synthetic DDD estimator. This transformed approach effectively addresses violations of parallel trends across both time and subgroup dimensions, as SC constructs a more accurate counterfactual based on the adjusted outcome. Thus, directly combining SC with DDD might be redundant when the transformed DDD approach already achieves the same goal.